# Faster shortest-path algorithms using the acyclic-connected tree

### v1

@@ 2.1 Graphs
In this paper, a graph refers to a weighted directed graph with a distinguished node called the source. Specifically, a graph is a tuple $G=(V,E,s)$ where $V$ is a set of nodes, $s\in V$ the source, and $E\subset V\times V$ a set of arcs. Most of our paper will use only this data, but to find shortest paths we will add a non-negative-valued weight function $\ell:E\to_{\geq 0}$ which stores the weight for each arc, with graph denoted by $G=(V,E,s,\ell)$.
We assume that all nodes in $G$ are reachable from $s$. This assumption is mild, because we could easily prune unreachable nodes prior to searching from $s$. Our goal is to solve the single-source shortest path problem.
@@
A module of $G=(V,E,s)$ is a set $M\subseteq V$ of nodes such that there exists an $m\in M$ where for every arc $u\xrightarrow{}v$, if $u\not\in M$ and $v\in M$ then $v=m$. In other words, all arcs into $M$ go to $m$. We call $m$ the source of $M$  [^2] Note that ‘modules’ also have a distinct interpretation in graph theory [22, 27]. Here we use a form specific to single-source graphs, which is closer to the usage in [10, 11]..
@@
Let $M$ and $H$ be modules with sources $m$ and $h$ respectively. Let $x\in M\cap H$. We know $m\geq_{s}x$ and $h\geq_{s}x$ so either (1) $m\geq_{s}h$ or (2) $h\geq_{s}m$ by Lemma 2.3. Assume w.l.o.g that (1) holds. Then any path $s\stackrel{{\scriptstyle}}{{\leadsto}}x$ contains first $m$, then $h$, then $x$. Because the subpath $h\stackrel{{\scriptstyle}}{{\leadsto}}x$ does not contain $m$, but $x\in M$, we must have $h\in M\cap H$. Claim: if $m\in M\cap H$ and $h\in M\cap H$ then $m=h$. To prove the claim, suppose that $h\in M\cap H$ and $m\in M\cap H$. Since there are paths $s\stackrel{{\scriptstyle}}{{\leadsto}}m$ and $s\stackrel{{\scriptstyle}}{{\leadsto}}h$, there must be an arc $u\xrightarrow{}v$ with $v\in M\cup H$ and $u\not\in M\cup H$ [^3] Unless $s\in M\cup H$. However, if $s\in M\cup H$, then $m=s$ or $h=s$ and since both $m$ and $h$ are in $M\cap H$, $s\in M\cap H$. Thus, $s$ must be the source of both $M$ and $H$ and the claim follows.. We know that $v=m$ or $v=h$, so $v\in M\cap H$. But then because $M$ and $H$ are modules, we must have $v=m=h$, completing the claim. If follows from this claim that either $m=h$, or $m\in M\setminus H$ and $h\in M\cap H$. Now let $u\xrightarrow{}v$ be such that $u\not\in M\cap H$ and $v\in M\cap H$. If $u\not\in M$, then $v=m$ as $M$ is a module, and so $m\in M\cap H$ implies $m=h$. If instead we have $u\not\in H$, this implies that $v=h$ as $H$ is a module. Hence $M\cap H$ is a module with source $h$. Similarly, let $u\xrightarrow{}v$ be such that $u\not\in M\cup H$ and $v\in M\cup H$. If $v\in M$, then $v=m$ as $M$ is a module. If $v\in H$, then $v=h$ as $H$ is a module. However, $h\in M\cap H$, so in this case $v\in M$ which implies that $v=m=h$. Hence $M\cup H$ is a module with source $m$. ∎
Finally, as discussed in the introduction, SSSP can be solved in linear time for DAGs. A simple linear algorithm for this is to first compute the topological order of the nodes and then visit the nodes in this order, updating distances to successor nodes as we visit each node [53]. This subroutine will play a key role in our algorithms since we work with nested acyclic graphs.
@@ 2.2 Nesting width
In Example 1.1, we could treat the set $S=\{r,d,e,f,g\}$ as a nested graph $G_{2}$ with source node $r$ precisely because all arcs into $S$ goes to $r$. In other words, we could represent this as a nested graph precisely because $S$ is a module. Note that if $S$ were not a module, say because an arc $b\xrightarrow{}d$ existed in Figure 1(a), then we would not be able to assign a unique source node to $G_{2}$. Hence the sets of modules define nesting decompositions of a graph.
@@
Note that every graph has at least one trivial nesting decomposition, consisting of only the trivial modules $V$ and $\{v\}$ for each $v\in V$. As a tree, this is a single root node with all singletons as its children. In particular, we always have the upper bound $\operatorname{nw}(G)\leq n$. Generally, though, graphs can have many distinct nesting decompositions. We want to find the ‘best’ nesting decomposition, whose width is minimal and thus equal to the nesting width.
@@ 3 The Acyclic-Connected Tree
Nesting decompositions are defined by modules, which are themselves related to the dominance relation $\geq_{s}$. This is the first ingredient of our decomposition: using the dominator tree to compute the relation $\geq_{s}$. However, dominance alone is not enough to construct a minimal-width nesting decomposition. Consider the following example: a graph with nodes $s,a_{i},b_{i}$ for $i$ from 1 to $N$, where there are arcs from $s$ to $a_{1}$ and $b_{1}$, and arcs from $a_{i}$ and $b_{i}$ to both $a_{i+1}$ and $b_{i+1}$. The dominator tree of this graph is the trivial one-layer tree with root $s$. However, this is acyclic, so a nesting decomposition of width $2$ exists (follows from Theorem 4.1). More generally, while the dominator tree tells us that all modules have source node $s$, this isn’t sufficient to identify them.
@@
Let $G=(V,E,s)$ be a graph, and $T$ be its dominator tree. For any node $a$ in $T$, let $D(a)$ denote its descendants and $C(a)$ its children. For each $a$, we define a graph $G_{a}$ whose node set is $C(a)$, with an arc $u\xrightarrow{}v$ in $G_{a}$ if $u\neq v$ and there is an arc from $D(u)$ to $D(v)$ in $G[D(a)]$ (the induced subgraph of $G$ consisting of descendants of $a$). The acyclic-connected tree (A-C tree) of $G$ is a map $a\mapsto S^{a}_{1},\dots,S^{a}_{k_{a}}$ taking each node $a$ to the sequence of strongly connected components of $G_{a}$ in topological order [^4] There may be many topological orders, so this is not unique. However, the choice of ordering will not affect any analysis, so we will treat it as unique for simplicity.. We depict this as in Figure 2.
@@
Figure 2: The graph $G$ from Example 1.1 (Fig. 2(a)) and its associated A-C tree (Fig. 2(b)), where the children of a node are topologically ordered from left to right. The child of each node $x$ in the tree is a set of nodes which represent a strongly connected component of $G_{x}$. We represent this by grouping the nodes in a rectangle (e.g., $e$ and $d$). We omit this rectangle for clarity if the strongly connected component is a singleton.
@@ 5 Computing the A-C tree
We proved in Theorem 4.1 that the A-C tree defines a nesting decomposition of minimal width. However, in order to use this to compute shortest paths we must also compute the A-C tree very efficiently. In this section we combine some classical linear-time algorithms (dominator trees, depth-first-search and strongly connected components) to achieve an algorithm which computes the A-C tree in linear time (Algorithm 1).
The first (and most technical) step of A-C tree computation is construction of the dominator tree. Fortunately, several decades of algorithmic work has established linear-time algorithms for this problem [4]. The second step, recalling Definition 3.1, is to group the children $C(a)$ of a node $a$ in the dominator tree into components, defined by reachability in the induced subgraph $D(a)$. The main challenge of this step is efficiently constructing the induced graphs $G_{a}$ whose nodes are $C(a)$ and where there is an arc $u\xrightarrow{}v$ in $G_{a}$ if there is a path $u\stackrel{{\scriptstyle}}{{\leadsto}}v$ in $G[D(a)]$. In Algorithm 2, we show how to construct the graphs $G_{a}$ for every $a$ in linear time using a variant of depth-first search. Once each graph $G_{a}$ has been constructed, we can run a linear-time strongly connected components algorithm (such as [53]) on each $G_{a}$ from $a$ to generate the components of $G_{a}$ in topological order, giving the children of $a$ in the A-C tree.
@@ 5       Set $T_{AC}[a]\leftarrow K_{1},K_{2},\dots,K_{m}$;
Output : $G_{p}$ for each node $p$ given by $graphs$
@@ 1 Run RecursiveDominanceGraphs ($s$);
2 $s[p]\leftarrow p.\textnormal{{graph}}$ graph for each node $p$;
3 return $graphs$;
@@ 10                   RecursiveDominanceGraphs $(w)$
Given a graph $G=(V,E,s)$, Algorithm 1 constructs its A-C tree in linear time $O(n+e)$.
@@
Algorithm 1 begins by constructing the dominator tree $T$ of $G$, which can be done in linear time [4]. Then we run the subroutine DominanceGraphs$(G,s,T)$ (Algorithm 2).
@@
[…]
@@ 6 Shortest paths from the A-C tree
As we showed in the introduction, the A-C tree allows us to construct a nesting decomposition of the graph, and from a nesting decomposition we can obtain a solution to SSSP by running a standard shortest path algorithm on each nested graph. This gives us an $O(e+n\log\operatorname{nw}(G))$ time algorithm. However, we can also achieve this running time in a much simpler and more direct way from the A-C tree, just by modifying a few lines of Dijkstra’s algorithm.
Our method, which we call Recursive Dijkstra’s, is described in Algorithm 3. The idea is simply that, upon visiting a node $a$, we first recursively search its children $K^{a}_{1},K^{a}_{2},\dots,K^{a}_{m}$, where again we search recursively when we reach new nodes. Each component $K^{a}_{i}$ contains its own priority queue, which are updated individually.
@@
Recursive Dijkstra’s solves SSSP in $O(e+n\log\operatorname{nw}(G))$ time.
@@ 22                         $K\leftarrow$ component $K^{a}_{i}$ containing $w$ in $T_{AC}$;
Algorithm 3 Recursive Dijkstra’s
@@
Overall, Recursive Dijkstra’s runs very similarly to the standard setup of Dijkstra’s algorithm, with two key differences: (1) for each node $a$ in the A-C tree it searches each connected component $K^{a}_{i}$ separately in topological order, and (2) it applies this process recursively using the A-C tree structure. In Algorithm 3, we highlight the differences from Dijkstra’s algorithm in blue. We will explain these below.
@@
Now we argue that this works in the recursive setting as well. Suppose that we are searching from a node $a$ in $T_{AC}$. Assume by induction that until this point, running Recursive Dijkstra’s on any node $x$ correctly identifies the shortest paths from $x$ to all nodes in $D(x)$. We will show that this holds as well for $a$. Let $K^{a}_{1},K^{a}_{2},\dots,K^{a}_{m}$ be the components of $a$, which again we will search in topological order. The only difference from the base case is that an arc $x\xrightarrow{}y$ in $G_{a}$ does not necessarily correspond to an arc $x\xrightarrow{}y$ in $G$—instead, it might correspond to an arc $u\xrightarrow{}w$ where $u\in D(x)$ and $w\in D(y)$. Suppose this is the case, for some $x$ and $y$ with $y\in K^{a}_{i}$. First, recall that because $D(y)$ is a module, $w=y$. Then, by the inductive hypothesis, we have already found the shortest path from $x$ to $u$, because $u\in D(x)$. Hence we can update the distance to $y$ in $K^{a}_{i}.queue$ by using $dist[u]+weight(u,y)$ (line 3). This operation takes place while searching recursively from $u$. That is, once we have run Recursive Dijkstra’s on $x$, we have correctly updated the distances to all successors $y$ in $G_{a}$. Hence searching the components in topological order in $G_{a}$ works as before, allowing us to correctly compute shortest paths from $a$ to all nodes in $D(a)$.
@@
Having proved the correctness, it remains to show the time complexity of Algorithm 3. To this end, note first that all subroutines in Algorithm 3 are done in linear time, except running Dijkstra’s algorithm (given by ComponentDijkstra) on each strongly connected component. We compute this time complexity. Totally analogous to Section 1.1, assuming $m$ strongly connected components $K_{1},\dots,K_{m}$ in the A-C tree, the time complexity to run Dijkstra’s algorithm on all these components is $O\big{(}(\sum_{i=1}^{m}e_{i})+(\sum_{i=1}^{m}n_{i})\log(\max_{i}n_{i})\big{)}$, where $n_{i}$ and $e_{i}$ is the number of nodes and edges in $K_{i}$. We note that $\sum_{i=1}^{m}n_{i}=n$ and $\sum_{i=1}^{m}e_{i}\leq e$, where the latter holds since some arcs may not be in the components. Moreover, $\max_{i}n_{i}=\max_{i}|K_{i}|=w_{AC}-1=\operatorname{nw}(G)-1\leq\operatorname{nw}(G)$ by the proof of Theorem 4.1. We conclude that the time complexity to run Dijkstra’s algorithm on all components is
$O\left(\left(\sum_{i=1}^{m}e_{i}\right)+\left(\sum_{i=1}^{m}n_{i}\right)\log\left(\max_{i}n_{i}\right)\right)\leq O\left(e+n\log\operatorname{nw}(G)\right).$
@@
This is also the time complexity of Algorithm 3 since all other subroutines are executed in linear time. ∎

### v2

@@ 2.1 Graphs
In this paper, a graph refers to a weighted directed graph with a distinguished node called the source. Specifically, a graph is a tuple $G=(V,E,s)$ where $V$ is a set of nodes, $s\in V$ the source, and $E\subset V\times V$ a set of arcs. Our graph decomposition uses only this data, but to find shortest paths we will also need a non-negative-valued weight function $\ell:E\to_{\geq 0}$ which stores the weight for each arc, with graph denoted by $G=(V,E,s,\ell)$.
We assume that all nodes in $G$ are reachable from $s$. This assumption is mild, because we could easily prune unreachable nodes prior to searching from $s$. Our goal is to solve the single-source shortest path problem:
@@
A module of $G=(V,E,s)$ is a set $M\subseteq V$ of nodes such that there exists an $m\in M$ where for every arc $u\xrightarrow{}v$, if $u\not\in M$ and $v\in M$ then $v=m$. In other words, all arcs into $M$ go to $m$. We call $m$ the source of MM. [^4] Note that ‘modules’ also have a distinct interpretation in graph theory [23, 28]. Here we use a form specific to single-source graphs, which is closer to the usage in [10, 11].
@@
See Appendix A. ∎
Finally, as discussed in the introduction, SSSP can be solved in linear time for DAGs, using the following simple method. First, one computes a topological ordering of the nodes and then visits the nodes in this order, updating distances to successor nodes as we visit each node [55]. This subroutine will play a key role in our algorithms since we work with nested acyclic graphs.
@@ 2.2 Nesting width
In Figure 1, we could treat the set $S$ as a nested graph $G_{2}$ with source node $h$ because all arcs into $S$ goes to $h$. In other words, we could represent $S$ as a nested graph precisely because it is a module. Note that if $S$ were not a module, say because an arc $b\xrightarrow{}d$ existed in Figure 1(a), then we would not be able to assign a unique source node to $G_{2}$. Hence the sets of modules define the nesting decompositions of a graph.
@@
Note that every graph has at least one trivial nesting decomposition, consisting of only the trivial modules $V$ and $\{v\}$ for each $v\in V$. As a tree, this is a single root node with all singletons as its children. In particular, we always have the upper bound $\operatorname{nw}(G)\leq n$. Generally, though, graphs can have many distinct nesting decompositions. We want to find the ‘best’ nesting decomposition, whose width is minimal and thus equal to the nesting width. Such a decomposition is given in the next section.
We finish this section by stating how the nesting width generalises DAGs given by the following proposition.
Proposition 2.9 (Nesting width and DAGs).
For any graph, [^5] For $G$ with more than one node. When $G$ has just one node, then $\operatorname{nw}(G)=1$. $2\leq\operatorname{nw}(G)\leq n$. In particular, for a DAG $G$, the lower bound $\operatorname{nw}(G)=2$ is attained. In this sense, $\operatorname{nw}(G)$ can be seen as a measure of how acyclic $G$ is.
Proof.
See Appendix A. ∎
Remark 1 (Nesting width and sparsity).
Sparse graphs are more likely to have small nesting width, but sparse graphs and bounded nesting width graphs are distinct concepts. For details, see Appendix D.
@@ 3 The Acyclic-Connected Tree
Nesting decompositions are defined by modules, which are themselves related to the dominance relation $\geq_{s}$. This is the first ingredient of our decomposition: using the dominator tree to compute the relation $\geq_{s}$. However, dominance alone is not enough to construct a minimal-width nesting decomposition. Consider the following example: a graph with nodes $s,a_{i},b_{i}$ for $i$ from 1 to $N$, where there are arcs from $s$ to $a_{1}$ and $b_{1}$, and arcs from $a_{i}$ and $b_{i}$ to both $a_{i+1}$ and $b_{i+1}$. The dominator tree of this graph is the trivial one-layer tree with root $s$. However, this is acyclic, so a nesting decomposition of width $2$ exists (follows from Proposition 2.9). More generally, while the dominator tree tells us that all modules have source node $s$, this is not sufficient to identify them.
@@
Let $G=(V,E,s)$ be a graph, and $T$ be its dominator tree. For any node $a$ in $T$, let $D(a)$ denote its descendants and $C(a)$ its children. For each $a$, we define a graph $G_{a}$ whose node set is $C(a)$, with an arc $u\xrightarrow{}v$ in $G_{a}$ if $u\neq v$ and there is an arc from $D(u)$ to $D(v)$ in $G[D(a)]$ (the induced subgraph of $G$ consisting of descendants of $a$). The acyclic-connected tree (A-C tree) of $G$ is a map $a\mapsto S^{a}_{1},\dots,S^{a}_{k_{a}}$ taking each node $a$ to the sequence of strongly connected components of $G_{a}$ in topological order. [^6] There may be many topological orders, so this is not unique. However, the choice of ordering will not affect any analysis, so we will treat it as unique for simplicity. We depict this as in Figure 2.
@@
Figure 2: The graph $G$ from Figure 1 and its associated A-C tree (Fig. 2(b)), where the children of a node are topologically ordered from left to right. The child of each node $x$ in the tree is a set of nodes which represent a strongly connected component of $G_{x}$. We represent this by grouping the nodes in a rectangle (e.g., $e$ and $d$). We omit this rectangle for clarity if the strongly connected component is a singleton.
@@ 5 Computing the A-C tree
We proved in Theorem 4.1 that the A-C tree defines a nesting decomposition of minimal width. However, for this to be useful for solving SSSP, we must also be able to compute the A-C tree efficiently. In this section we combine some classical linear-time algorithms (dominator trees, depth-first-search and strongly connected components) to achieve a linear-time algorithm for computing the A-C tree (Algorithm 1).
The first (and most technical) step of A-C tree computation is the construction of the dominator tree. Fortunately, several decades of algorithmic work has established linear-time algorithms for this problem [4]. The second step, recalling Definition 3.1, is to group the children $C(a)$ of a node $a$ in the dominator tree into components, defined by reachability in the induced subgraph $D(a)$. The main challenge of this step is to efficiently construct the induced graphs $G_{a}$ whose nodes are $C(a)$ and where there is an arc $u\xrightarrow{}v$ in $G_{a}$ if there is a path $u\stackrel{{\scriptstyle}}{{\leadsto}}v$ in $G[D(a)]$. In Algorithm 2, we show how to construct the graphs $G_{a}$ for every $a$ in linear time using a variant of depth-first search. Once each graph $G_{a}$ has been constructed, we can run a linear-time strongly connected components algorithm (such as [55]) on each $G_{a}$ from $a$ to generate the components of $G_{a}$ in topological order, giving the children of $a$ in the A-C tree.
@@ 5   Set $T_{AC}[a]\leftarrow K_{1},K_{2},\dots,K_{m}$;
Output : $G_{p}$ for each node $p$ given by graphsgraphs
@@ 1 Run RecursiveDominanceGraphs ($s$);
2 $graphs[p]\leftarrow p.\textnormal{{graph}}$ for each node $p$;
3 return graphsgraphs;
@@ 10       RecursiveDominanceGraphs $(w)$
Given a graph $G=(V,E,s)$, Algorithm 1 constructs its A-C tree in time $O(n+m)$.
@@
Algorithm 1 begins by constructing the dominator tree $T$ of $G$, which can be done in linear time [4]. Then we run the subroutine DominanceGraphs(G,s,T)(G,s,T) (Algorithm 2).
@@
Fix a node $a$. We would like to show that $a.graph$ equals $G_{a}$ after running Algorithm 2. To do this, we must show that for every distinct $x,y\in C(a)$ where there is an arc from $D(x)$ to $D(y)$ in $G[D(a)]$, we add the arc $x\xrightarrow{}y$ to $a.graph$, and conversely, every added arc $x\xrightarrow{}y$ to $a.graph$ must be an on this form. First, observe that at the time we call $\texttt{RecursiveDominanceGraphs}(a)$, no descendants of $a$ have yet been visited. By running a DFS from $a$ in the dominator tree, $\texttt{RecursiveDominanceGraphs}(a)$ visits every descendant of $a$ exactly once. For each $v\in D(a)$, we examine each outgoing arc $v\xrightarrow{}w$ exactly once.
To this end, suppose first there is an arc $v\xrightarrow{}w$ in $G[D(a)]$ where $v\in D(x)$ and $w\in D(y)$ for distinct $x,y\in C(a)$. Then, while running $\texttt{RecursiveDominanceGraphs}(x)$, we reach the node $v$ and examine the arc $v\xrightarrow{}w$. Because $T$ is a dominator tree, $D(y)$ is a module with source $y$. Thus, any arc into $D(y)$ from outside $D(y)$ must go to $y$. Since $v\notin D(y)$, the arc $v\xrightarrow{}w$ is such an arc, and therefore $w=y$. Then $w.parent=y.parent=a$. Because we are currently running $\texttt{RecursiveDominanceGraphs}(x)$, $a.current=x$. Hence on line 2 we add the arc $x\xrightarrow{}y$ to $a.graph$.
@@ 6 Shortest paths from the A-C tree
From the A-C tree, we implicitly obtain a minimum-width nesting decomposition. In this section, we show how this can be exploited to improve state-of-the-art shortest-path algorithms.
First (Section 6.1), we show how, using the A-C tree, we can modify just a few lines of Dijkstra’s algorithm and obtain an algorithm (Algorithm 3) with improved time complexity $O(m+n\log\operatorname{nw}(G))$. To the best of our knowledge, this is currently the best asymptotic improvement of Dijkstra’s algorithm for general graphs. In particular, for classes of graphs with bounded nesting width, Algorithm 3 runs in linear time.
Second (Section 6.2), we show how to apply the A-C tree to any any black-box SSSP algorithm (such as the recent work of [19]). More precisely, we present an algorithm called Recursive SSSP (Algorithm 4) that takes as input a black-box SSSP algorithm $\mathcal{SP}$ and applies it recursively using the A-C tree. Given a time complexity bound $f_{\mathcal{SP}}(n,m)$ for $\mathcal{SP}$, the time complexity of Recursive SSSP is $O(m\alpha(n)+\sum_{K_{i}}f_{\mathcal{SP}}(n_{i},m_{i}))$ (Theorem 6.3), where the sum is taken over all components $K_{i}$ of the A-C tree, $n_{i}$ ($m_{i}$) is the number of internal nodes (arcs) in component $K_{i}$, and $\alpha(n)$ is the extremely slow-growing inverse Ackermann function. This results in improved time complexity under mild conditions. Indeed, whenever $f_{\mathcal{SP}}(n,m)\in\Omega(m\alpha(n))$ this is an asymptotic improvement. To illustrate this, we apply Recursive SSSP on the recent sparse graph algorithm of [19], which runs in time $O(m\log^{2/3}(n))$. The result is an algorithm with improved time complexity $O(m\alpha(n)+m\log^{2/3}(\operatorname{nw}(G)))$ (Corollary 6.4). In particular, for sparse graphs $m=O(n)$, our algorithm runs in $O(n\alpha(n)+n\log^{2/3}(\operatorname{nw}(G)))$. To the best of our knowledge, this is best beyond-worst-case complexity bound for SSSP on sparse graphs.
6.1 Recursive Dijkstra
We will begin with Algorithm 3, which we call Recursive Dijkstra. The idea is simply that, upon visiting a node $a$, we first recursively search its children $K^{a}_{1},K^{a}_{2},\dots,K^{a}_{m}$ in the A-C tree, where again we search recursively when we reach new nodes. Each component $K^{a}_{i}$ contains its own priority queue, which is updated individually.
@@ 6.1 Recursive Dijkstra
Recursive Dijkstra solves SSSP in $O(m+n\log\operatorname{nw}(G))$ time.
@@ 22           $K\leftarrow$ component $K^{a}_{i}$ containing $w$ in $T_{AC}$;
Algorithm 3 Recursive Dijkstra
@@
Correctness. Overall, Recursive Dijkstra runs very similarly to the standard setup of Dijkstra’s algorithm, with two key differences: (1) for each node $a$ in the A-C tree it searches each connected component $K^{a}_{i}$ separately in topological order, and (2) it applies this process recursively using the A-C tree structure. In Algorithm 3, we highlight the differences from Dijkstra’s algorithm in blue. We will explain these below.
@@
Now we argue that this works in the recursive setting as well. Suppose that we are searching from a node $a$ in $T_{AC}$. Assume by induction that until this point, running Recursive Dijkstra on any node $x$ correctly identifies the shortest paths from $x$ to all nodes in $D(x)$. We will show that this holds as well for $a$. Let $K^{a}_{1},K^{a}_{2},\dots,K^{a}_{m}$ be the components of $a$, which again we will search in topological order. The only difference from the base case is that an arc $x\xrightarrow{}y$ in $G_{a}$ does not necessarily correspond to an arc $x\xrightarrow{}y$ in GG—instead, it might correspond to an arc $u\xrightarrow{}w$ where $u\in D(x)$ and $w\in D(y)$. Suppose this is the case, for some $x$ and $y$ with $y\in K^{a}_{i}$. First, recall that because $D(y)$ is a module, $w=y$. Then, by the inductive hypothesis, we have already found the shortest path from $x$ to $u$, because $u\in D(x)$. Hence we can update the distance to $y$ in $K^{a}_{i}.queue$ by using $dist[u]+weight(u,y)$ (line 3). This operation takes place while searching recursively from $u$. That is, once we have run Recursive Dijkstra on $x$, we have correctly updated the distances to all successors $y$ in $G_{a}$. Hence searching the components in topological order in $G_{a}$ works as before, allowing us to correctly compute shortest paths from $a$ to all nodes in $D(a)$.
@@
Time complexity. Having proved the correctness, it remains to show the time complexity of Algorithm 3. To this end, note first that all subroutines in Algorithm 3 are done in linear time, except running Dijkstra’s algorithm (given by ComponentDijkstra) on each strongly connected component. We compute this time complexity. Totally analogous to Section 1.1, assuming $m$ strongly connected components $K_{1},\dots,K_{m}$ in the A-C tree, the time complexity to run Dijkstra’s algorithm on all these components is $O\big((\sum_{i=1}^{m}m_{i})+(\sum_{i=1}^{m}n_{i})\log(\max_{i}n_{i})\big)$, where $n_{i}$ and $m_{i}$ is the number of nodes and edges in $K_{i}$. We note that $\sum_{i=1}^{m}n_{i}=n$ and $\sum_{i=1}^{m}m_{i}\leq m$, where the latter holds since some arcs may not be in the components. Moreover, $\max_{i}n_{i}=\max_{i}|K_{i}|=w_{AC}-1=\operatorname{nw}(G)-1\leq\operatorname{nw}(G)$ by the proof of Theorem 4.1. We conclude that the time complexity to run Dijkstra’s algorithm on all components is
$O\left(\left(\sum_{i=1}^{m}m_{i}\right)+\left(\sum_{i=1}^{m}n_{i}\right)\log\left(\max_{i}n_{i}\right)\right)\leq O\left(m+n\log\operatorname{nw}(G)\right).$
@@
This is also the time complexity of Algorithm 3, as all other subroutines are executed in linear time. ∎
Theorem 6.1 shows that Recursive Dijkstra is asymptotically faster than Dijkstra’s algorithm (since $\operatorname{nw}(G)\leq n$). In particular, for graphs of bounded nesting width, Recursive Dijkstra runs in linear time:
Corollary 6.2.
If a class of graphs has bounded nesting width, then Recursive Dijkstra runs in linear time $O(n+m)$ on this class.
Input : SSSP algorithm $\mathcal{SP}$ and graph $G=(V,E,s)$
Output : Shortest distances and path tree given by dist and parent
1 Create dictionaries dist, parent, originorigin;
2 Initialise $\textnormal{{dist}}[s]=0$, and $\textnormal{{dist}}[v]=\infty$ and $\textnormal{{parent}}[v]\leftarrow None$ for all $v$;
3 Initialise $v.find=v$ and $v.find.weight=0$ for each node $v$ in $G$;
4 Compute $T_{AC}\leftarrow\texttt{A-C\_tree}(G)$;
5 TreeRecursion $(s)$;
6 Compute dist using the shortest path tree parent ;
7 return dist, parent ;
8 TreeRecursion $(a)$
9 $K_{1},K_{2},\dots,K_{m}\leftarrow T_{AC}[a]$;
10 for each $i$ from 1 to $m$ do
11      Create empty graph $G_{K_{i}}$, with node set $K_{i}\cup\{a\}$;
12    for each $u\in K_{i}$ do
13       TreeRecursion ($u$);
14
15    for each $u\in K_{i}$ do
16       for each arc $(x,u)$ in $G$ with weight $W$ do
17          $v,d\leftarrow find(x)$;
18          if $(v,u)$ in $G_{K_{i}}$ then
19             if $d+W<weight_{G_{K_{i}}}(v,u)$ then
20                  Update weight of $(v,u)$ in $G_{K_{i}}$ to $d+W$;
21                $origin[v,u]\leftarrow x$;
22
23          else
24               Add arc $(v,u)$ to $G_{K_{i}}$ with weight $d+W$;
25             $origin[v,u]\leftarrow x$;
26
    $dist\_a\_K_{i},\ parent\_a\_K_{i}\leftarrow\mathcal{SP}(G_{K_{i}},a)$;
    // SP:s to $K_{i}$ from $a$
27    for each $u\in K_{i}$ do
28       $(u.find,u.find.weight)\leftarrow(a,dist\_a\_K_{i}[u])$ ;
29       $\textnormal{{parent}}[u]\leftarrow origin[parent\_a\_K_{i}[u],u]$
Algorithm 4 Recursive SSSP
6.2 Recursive SSSP
In Algorithm 3, we integrated the A-C tree into the recursive structure of Dijkstra’s algorithm. Here, we show that the A-C tree can be used to speed up any SSSP algorithm in a black-box way, using a slightly more complicated method shown in Algorithm 4. It operates on the same principles as Recursive Dijkstra, in that we will recursively find shortest paths in each subtree of the A-C tree, except that we can no longer assume that when we recursively find the shortest paths from a node $u$, the shortest path from $s$ to $u$ is already known. To account for this, we use (1) an augmented merge-find data structure (called at line 4 using findfind) to hold the roots of recursive shortest-path-subtrees, see Appendix C for details, and (2) build an auxiliary graph $G_{K_{i}}$ for each component $K_{i}$ in the A-C tree on which we will then separately run the SSSP algorithm (line 4). Because we run SSSP on a modified graph, we must convert the shortest paths found in $G_{K_{i}}$ back into paths in $G$. We do this by storing the ‘true’ tail of an arc $v\xrightarrow{}u$ in $G_{K_{i}}$ in a dictionary $origin[v,u]$, from which we recover the true shortest-path-parents on line 4. The formal result is given in the following theorem. We defer the proof to Appendix B, though the principles are the same as for Recursive Dijkstra.
Theorem 6.3.
Given an SSSP algorithm $\mathcal{SP}$ that runs in time $f_{\mathcal{SP}}(n,m)$ on a graph with $n$ nodes and $m$ arcs, Algorithm 4 solves SSSP in time $O(m\alpha(n)+\sum_{K_{i}}f_{\mathcal{SP}}(n_{i},m_{i}))$, where the sum is taken over all components $K_{i}$ of the A-C tree, $n_{i}$ ($m_{i}$) is the number of internal nodes (respectively arcs [^9] Specifically, an arc is ‘internal’ in $K_{i}$ with respect to $T[a]$ if it corresponds to an arc between two nodes of $K_{i}$ in $G_{a}$.) in component $K_{i}$, and $\alpha(n)$ is the inverse Ackermann function.
Proof.
See Appendix B. ∎
Corollary 6.4.
Applying Algorithm 4 to the SSSP algorithm in [19] (with time complexity $O(m\log^{2/3}(n)$), results in an SSSP algorithm with improved time complexity $O(m\alpha(n)+m\log^{2/3}(\operatorname{nw}(G)))$.
Proof.
See Appendix B. ∎